### v2

@@ 1 Introduction
The analysis of multiagent networks has been applied to a variety of subjects, such as multi-robot systems [17], wireless sensor networks [11], and large language models [16]. One of the simplest and most fundamental algorithms in this context is linear consensus, also known as average consensus. The linear consensus algorithm is used when each agent has a scalar value and have to update its value toward the weighted average of initial values, relying only on limited and decentralized communication. Despite its simplicity, the linear consensus algorithm forms the foundation for many tasks in network coordination, such as formation control [22, 27], distributed optimization [31], and cooperative leader following [18, 30].
@@
The performance measurement of the linear consensus algorithm has been studied in various ways. A linear consensus network can be understood as a linear discrete-time dynamical system closely related to a Markov chain [5], making Markov chain theory a key analytical tool. In fact, the performance of the linear consensus algorithm has been analyzed by the speed of convergence using spectral analysis of the transition matrix of the associated Markov chain [4, 29]. On the other hand, recent studies often adopt approaches from a control-theoretic point of view, such as linear quadratic (LQ) cost, which is classical in control. The LQ cost is defined as the sum of squared $L_{2}$-norm of the difference between the state trajectory and the final state. Moreover, the LQ cost appears in the error estimation of noisy consensus, a variant of linear consensus with additive random noise. Motivated by these connections, several studies have employed the LQ cost or similar functions as performance measures [5, 14, 21]. In these studies, various approaches are employed:
@@
  • Eigenvalue calculation: Eigenvalues are calculated for special graph types, including Cayley graphs, grid graphs, and random geometric graphs [5, 14].
@@
  • Effective resistance: The concept, which bridges electrical networks with Markov chains [10], has been employed in several recent studies [24, 26]. Notably, previous research has estimated the LQ cost using the effective resistance [21], enabling the assessment of LQ cost based on the network’s graph topology.
@@
  • Hitting time of a Markov chain (see [1]): Recent research [19] has provided an exact formulation of a variant of the LQ cost by leveraging hitting times.
@@
In this paper, we derive both upper and lower bounds on the LQ cost for nonreversible cases using effective resistance. This generalizes the estimation of the LQ cost for reversible linear consensus networks, as studied in [21], to nonreversible networks. A key advantage of nonreversible Markov chains is its applicability to directed networks. While the reversibility is a useful assumption, it does not hold in directed communication scenarios, making the nonreversible setting highly relevant. Thus, recent studies have highlighted the importance of nonreversible Markov chains [6, 28, 7]. A central technique in this line is called reversiblization of the transition matrix, which allows one to exploit results of studies on reversible Markov chains [13].
@@
    We relate the reversiblization and the original nonreversible Markov chain by introducing new notions named the back-and-forth path and pivot node. These concepts serve as alternatives to the notion of $2$-fuzz of an undirected graph [2], which is central in reversible cases, but cannot directly apply to nonreversible cases. These concepts, mainly used in Lemma 3.16, demonstrate how graph-theoretic methods can be adapted to the analysis of nonreversible cases.
@@
Furthermore, to clarify the behavior of our bounds, we introduce some specific applications and experiments, including Cayley graphs and random geometric graphs. Cayley graphs are highly symmetric and thus analytically tractable, while random geometric graphs model real-world communication networks. These graphs have been extensively studied in the literature (see [12] and [5]).
@@
The remainder of this paper is organized as follows: In Section 2, we introduce the mathematical formulation of linear consensus algorithm, effective resistance, and the relationship between them. In Section 3, we present our main result on the estimation of the LQ cost, along with its proofs. Section 4 demonstrates applications and and numerical experiments, and Section 5 concludes the paper.
@@
Let $\mathbb{R}$ be the set of real numbers, $\mathbb{R}^{n}$ be the set of $n$-dimensional real vectors, and $\mathbb{R}^{n\times m}$ be the set of $n\times m$ real matrices. We denote the $i$-th entry of $\bm{v}\in\mathbb{R}^{n}$ by $v_{i}$, and the $(i,j)$ entry of $A\in\mathbb{R}^{n\times m}$ by $A_{ij}$. Let $\bm{e}_{i}\in\mathbb{R}^{n}$ be the vector whose $i$-th entry is $1$ and other entries are 0, and $\bm{1}\in\mathbb{R}^{n}$ be the vector with all entries equal to $1$. We denote the identity matrix by $I$ and zero matrix by $O$. For a vector $\bm{v}\in\mathbb{R}^{n}$, let $\bm{v}^{\mathsf{T}}$ be the transpose of $\bm{v}$, $\operatorname{span}\{\bm{v}\}$ be a vector space spanned by $\bm{v}$, and $\operatorname{diag}(\bm{v})\in\mathbb{R}^{n\times n}$ be the diagonal matrix satisfying $\operatorname{diag}(\bm{v})_{ii}=v_{i}$ for all $i=1,2,\dots,n$ and $\operatorname{diag}(\bm{v})_{ij}=0$ if $i\neq j$. For a matrix $A$, let $A^{\mathsf{T}}$ be the transpose of $A$, $\operatorname{tr}A$ be the trace of $A$, $\ker A$ be the kernel of $A$, and $A^{\dagger}$ be the Moore-Penrose pseudoinverse of $A$. For two matrices $A$ and $B$, $A\geq B$ means $A_{ij}\geq B_{ij}$ for all $i$ and $j$. We denote a graph with a set of nodes $V$ and a set of edges $\mathcal{E}$ by $(V,\mathcal{E})$. For an undirected graph, we denote an edge between nodes $u$ and $v$ by $\{u,v\}$, while for a directed graph, we denote an edge from node $u$ to node $v$ by $(u,v)$.
@@
The reversiblization of the consensus matrix $P$
@@ 2.1 Definition of linear consensus algorithm
Throughout this paper, we assume two properties: $P$ is irreducible, and the diagonal elements of $P$ are positive, where $P$ is irreducible if and only if $\mathcal{G}_{\mathrm{dir}}(P)$ is strongly connected. The positive diagonal assumption may appear to be strong, but it is natural for each agent to use its current state for each iteration, and easy to implement since it does not require any additional communication. In addition, this assumption leads to some useful properties, such as aperiodicity of $P$, which means that the greatest common divisor of the lengths of all cycles in $\mathcal{G}_{\mathrm{dir}}(P)$ is $1$. Aperiodicity is a nice property, it is usual to ensure this property by adding some assumptions. For example, the PageRank algorithm calculates the stationary distribution of a Google matrix, which is arranged from the initial hyperlink matrix to make it aperiodic [20].
By these two assumptions, Perron-Frobenius theorem [25] implies that $P$ has the eigenvalue $1$ with the multiplicity $1$, the corresponding right eigenvector is $\bm{1}$, and corresponding left eigenvector is a strictly positive vector. We denote the left eigenvector by $\bm{\pi}^{\mathsf{T}}$ normalized so that $\sum_{u}\pi_{u}=1$, and call $\bm{\pi}$ the invariant measure of $P$. It also holds that, when $t$ grows to infinity, $P^{t}$ converges to $\bm{1}\bm{\pi}^{\mathsf{T}}$. Because of (1), we obtain $\lim_{t\to\infty}\bm{x}(t)=\bm{1}\bm{\pi}^{\mathsf{T}}\bm{x}(0)$. We also use the diagonal matrix $\varPi:=\operatorname{diag}(\bm{\pi})$.
@@ 2.2 Performance measure of linear consensus algorithm
As a performance measure, we use the linear quadratic (LQ) cost [21]:
@@
The cost (2) is obtained by evaluating $\mathrm{E}\left[\sum_{t\geq 0}\|\bm{x}(t)-\bm{x}(\infty)\|^{2}\right]$, where $\mathrm{E}[\cdot]$ means the expected value, under the assumption that $\bm{x}(0)$ is a random vector with covariance $\mathrm{E}\left[\bm{x}(0)\bm{x}(0)^{\mathsf{T}}\right]=I$ [14, 21]. In addition, the cost (2) also appears in the noisy consensus [14, 21, 19]. The noisy consensus repeats the following update:
@@
(3) $\frac{1}{n}\lim_{t\to\infty}\mathrm{E}\left[\|\tilde{\bm{e}}\|^{2}\right]$ $=$ $J(P).$
From the context of noisy consensus, we also use another variant [19]:
@@
which is obtained by substituting $\|\tilde{\bm{e}}\|^{2}$ in (3) by $\sum_{i}\pi_{i}\tilde{e}_{i}^{2}$. Unlike (2), which sums up the errors uniformly, (4) sums up the errors with weights according to $\bm{\pi}$.
@@ 2.3 Effective resistance
We consider a voltage vector of the nodes $\bm{v}$, where $v_{a}$ represents the voltage of the node $a$, and a current vector $\bm{i}$, where $i_{a}$ represents the current flowing out of (or into when negative) the node $a$. By Ohm’s law, the current from $a$ to $b$ through an edge $\{a,b\}$ is $C_{ab}(v_{a}-v_{b})$. Therefore, by Kirchhoff’s law, $i_{a}$ is determined by:
@@
Because the current injection and extraction are balanced, we can assume $\bm{i}^{\mathsf{T}}\bm{1}=0$. Under this condition, the following lemma is well known fact (e.g., Lemma 6.12 in [3]).
@@ 2.4 The relationships between linear consensus algorithm and effective resistance
The proof is analogous to the argument of Section 3.1.2 in [21].
Let $\Phi(P):=\Phi_{n}(P)=n\varPi P$, where $n$ denotes the number of rows of $P$. Then we can see a clear relationship between the random walk with the transition matrix $P$ and some properties in the electrical network with conductance matrix $\Phi(P)$ such as the effective resistance, the voltage and the current (see [10]). In the analogy of random walk, a matrix called Green matrix [21] plays an important role. The Green matrix of consensus matrix $P$ is defined as:
@@
This matrix is also called the fundamental matrix [1]. For a consensus matrix $P$, this matrix is well defined (see Chapter 11.5 in [15]).
@@
Because $P^{t}-\bm{1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\pi}^{\mathsf{T}}}=(P-\bm{1}\bm{\pi})^{t}$ for $t\geq 1$, (7) can be reformulated as follows:
@@
A fundamental matrix often refers not to $G(P)$, but to the right-hand side of (8) or (9) (e.g. [15, 19]).
@@
which can be found in [21]. Such properties in this section are used to relate $J(P)$ and the effective resistance of the graph. In particular, $G(P)$ plays an important role in evaluating $J(P)$.
@@ 3 Performance analysis of linear consensus algorithm using effective resistance
In this section, we give upper and lower bounds for $J(P)$ using the concept of effective resistance. A bound for $J(P)$ has been known for reversible $P$ [21, 19], but not for nonreversible cases. To provide bounds for nonreversible $P$, we use a reversible matrix $P^{*}P$, where $P^{*}:=\varPi^{-1}P^{\mathsf{T}}\varPi$. Our result is a generalization of [21].
@@ 3.1 Properties of $P^{*}P$
We show some properties of $P^{*}P$. For every consensus matrix $P$, $P^{*}P$ is a reversible consensus matrix with the same invariant measure as $P$, so it is called a multiple reversiblization of $P$ [13]. When $P$ is a reversible consensus matrix, we have $P^{*}=\varPi^{-1}(\varPi P)^{\mathsf{T}}=\varPi^{-1}(\varPi P)=P$ and $P^{*}P=P^{2}$.
@@
Let $Q:=\varPi^{1/2}P\varPi^{-1/2}$. Then $\operatorname{tr}\left((P^{*})^{t}P^{t}\right)=\operatorname{tr}\left((Q^{t})^{\mathsf{T}}Q^{t}\right)=\sum_{i}\sigma_{i}^{2}(Q^{t})$, where $\sigma_{i}(Q)$ is the $i$-th singular value of $Q$. Moreover, $\operatorname{tr}\left((P^{*}P)^{t}\right)=\operatorname{tr}\left((Q^{\mathsf{T}}Q)^{t}\right)=\sum_{i}\sigma_{i}^{2t}(Q)$, because $(Q^{\mathsf{T}}Q)^{t}$ is symmetric. Using $\sigma_{i}(Q^{t})\leq\sigma_{i}^{t}(Q)$, as shown in 9.H.2.a of [23], we obtain (10).
The equality of (10) holds not only for reversible matrices. For example, normal consensus matrices satisfy the property $P^{*}P=PP^{*}$. The class of normal consensus matrices includes Cayley graphs, and $J(P)$ on Cayley graphs have been studied (e.g. [5]).
@@
Because $P$ is normal, $(I-P)(I-P^{\mathsf{T}})=(I-P^{\mathsf{T}})(I-P)$. Multiplying $\bm{1}$ from the right, we obtain $(I-P)(I-P^{\mathsf{T}})\bm{1}=\bm{0}$, which means $(I-P^{\mathsf{T}})\bm{1}\in\ker(I-P)$. Therefore, there exists some real value $a$ that satisfies $(I-P^{\mathsf{T}})\bm{1}=a\bm{1}$, because $\ker(I-P)=\operatorname{span}\{\bm{1}\}$. Then we multiply $\bm{1}^{\mathsf{T}}$ from the left and obtain $\bm{0}=an$. This means that $a=0$ and $(I-P^{\mathsf{T}})\bm{1}=\bm{0}$. Thus, $P$ is doubly-stochastic and the invariant measure $\bm{\pi}$ is $\frac{1}{n}\bm{1}$. Therefore, we obtain $P^{*}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}P^{\mathsf{T}}}$ and $P^{*}P=P^{\mathsf{T}}P=PP^{\mathsf{T}}=PP^{*}$.

### v3

@@ 1 Introduction
The analysis of multiagent networks has been applied to a variety of subjects, such as multi-robot systems [robot], wireless sensor networks [WSN], and large language models [LLM]. One of the simplest and most fundamental algorithms in this context is linear consensus, also known as average consensus. The linear consensus algorithm is used when each agent has a scalar value and has to update its value toward the weighted average of initial values, relying only on limited and decentralized communication. Despite its simplicity, the linear consensus algorithm forms the foundation for many tasks in network coordination, such as formation control [formation1, multivehicle2], distributed optimization [optim1], and cooperative leader following [leader0, leader1].
@@
The performance measurement of the linear consensus algorithm has been studied in various ways. A linear consensus network can be understood as a linear discrete-time dynamical system closely related to a Markov chain [cayley], making Markov chain theory a key analytical tool. In fact, the performance of the linear consensus algorithm has been analyzed by the speed of convergence using spectral analysis of the transition matrix of the associated Markov chain [spectral1, spectral2]. On the other hand, recent studies often adopt approaches from a control-theoretic point of view, such as linear quadratic (LQ) cost, which is classical in control. The LQ cost is defined as the sum of squared $L_{2}$-norm of the difference between the state trajectory and the final state. Moreover, the LQ cost appears in the error estimation of noisy consensus, a variant of linear consensus with additive random noise. Motivated by these connections, several studies have employed the LQ cost or similar functions as performance measures [cayley, aboutlqcost, mainp]. In these studies, various approaches are employed:
@@
  • Eigenvalue calculation: Eigenvalues are calculated for special graph types, including Cayley graphs, grid graphs, and random geometric graphs [cayley, aboutlqcost].
@@
  • Effective resistance: The concept, which bridges electrical networks with Markov chains [randomwalks], has been employed in several recent studies [citeres1, satokazu]. Notably, previous research has estimated the LQ cost using the effective resistance [mainp], enabling the assessment of LQ cost based on the network’s graph topology.
@@
  • Hitting time of a Markov chain (see [revmarkov]): Recent research [novelnoisy] has provided an exact formulation of a variant of the LQ cost by leveraging hitting times.
@@
In this paper, we derive both upper and lower bounds on the LQ cost for nonreversible cases using effective resistance. This generalizes the estimation of the LQ cost for reversible linear consensus networks, as studied in [mainp], to nonreversible networks. A key advantage of nonreversible Markov chains is its applicability to directed networks. While the reversibility is a useful assumption, it does not hold in directed communication scenarios, making the nonreversible setting highly relevant. Thus, recent studies have highlighted the importance of nonreversible Markov chains [nonrev1, nonrev2, nonrev3]. A central technique in this line is called reversiblization of the transition matrix, which allows one to exploit results of studies on reversible Markov chains [reversiblization].
@@
    We relate the reversiblization and the original nonreversible Markov chain by introducing new notions named the back-and-forth path and pivot node. These concepts serve as alternatives to the notion of $2$-fuzz of an undirected graph [hfuzz], which is central in reversible cases, but cannot directly apply to nonreversible cases. These concepts, mainly used in Lemma 3.16, demonstrate how graph-theoretic methods can be adapted to the analysis of nonreversible cases.
@@
Furthermore, to clarify the behavior of our bounds, we introduce some specific applications and experiments, including Cayley graphs and random geometric graphs. Cayley graphs are highly symmetric and thus analytically tractable, while random geometric graphs model real-world communication networks. These graphs have been extensively studied in the literature (see [rggs] and [cayley]).
@@
The remainder of this paper is organized as follows: In Section 2, we introduce the mathematical formulation of linear consensus algorithm, effective resistance, and the relationship between them. In Section 3, we present our main result on the estimation of the LQ cost, along with its proofs. Section 4 demonstrates applications and numerical experiments, and Section 5 concludes the paper.
@@
Let $\mathbb{R}$ be the set of real numbers, $\mathbb{R}^{n}$ be the set of $n$-dimensional real vectors, and $\mathbb{R}^{n\times m}$ be the set of $n\times m$ real matrices. We denote the $i$-th entry of $\bm{v}\in\mathbb{R}^{n}$ by $v_{i}$, and the $(i,j)$ entry of $A\in\mathbb{R}^{n\times m}$ by $A_{ij}$. Let $\bm{e}_{i}\in\mathbb{R}^{n}$ be the vector whose $i$-th entry is $1$ and other entries are 0, and $\bm{1}\in\mathbb{R}^{n}$ be the vector with all entries equal to $1$. We denote the identity matrix by $I$ and the zero matrix by $O$. For a vector $\bm{v}\in\mathbb{R}^{n}$, let $\bm{v}^{\mathsf{T}}$ be the transpose of $\bm{v}$, $\operatorname{span}\{\bm{v}\}$ be a vector space spanned by $\bm{v}$, and $\operatorname{diag}(\bm{v})\in\mathbb{R}^{n\times n}$ be the diagonal matrix satisfying $\operatorname{diag}(\bm{v})_{ii}=v_{i}$ for all $i=1,2,\dots,n$ and $\operatorname{diag}(\bm{v})_{ij}=0$ if $i\neq j$. For a matrix $A$, let $A^{\mathsf{T}}$ be the transpose of $A$, $\operatorname{tr}A$ be the trace of $A$, $\ker A$ be the kernel of $A$, and $A^{\dagger}$ be the Moore-Penrose pseudoinverse of $A$. For two matrices $A$ and $B$, $A\geq B$ means $A_{ij}\geq B_{ij}$ for all $i$ and $j$. We denote a graph with a set of nodes $V$ and a set of edges $\mathcal{E}$ by $(V,\mathcal{E})$. For an undirected graph, we denote an edge between nodes $u$ and $v$ by $\{u,v\}$, while for a directed graph, we denote an edge from node $u$ to node $v$ by $(u,v)$.
@@
$\bm{\pi}$
The invariant measure of $P$
$\varPi$
The diagonal matrix defined as $\operatorname{diag}(\bm{\pi})$
@@
The matrix defined as $\varPi^{-1}P^{\mathsf{T}}\varPi$
@@ 2.1 Definition of linear consensus algorithm
Throughout this paper, we assume two properties: $P$ is irreducible, and the diagonal elements of $P$ are positive, where $P$ is irreducible if and only if $\mathcal{G}_{\mathrm{dir}}(P)$ is strongly connected. The positive diagonal assumption may appear to be strong, but it is natural for each agent to use its current state for each iteration, and easy to implement since it does not require any additional communication. In addition, this assumption leads to some useful properties, such as aperiodicity of $P$, which means that the greatest common divisor of the lengths of all cycles in $\mathcal{G}_{\mathrm{dir}}(P)$ is $1$. Aperiodicity is a desirable property and is often enforced by imposing additional assumptions. For example, the PageRank algorithm calculates the stationary distribution of a Google matrix, which is arranged from the initial hyperlink matrix to make it aperiodic [pagerank].
Under these two assumptions, the Perron-Frobenius theorem [perronfrobenius] implies that $P$ has the eigenvalue $1$ with the multiplicity $1$, the corresponding right eigenvector is $\bm{1}$, and corresponding left eigenvector is a strictly positive vector. We denote the left eigenvector by $\bm{\pi}^{\mathsf{T}}$ normalized so that $\sum_{u}\pi_{u}=1$, and call $\bm{\pi}$ the invariant measure of $P$. Moreover, as $t\to\infty$, $P^{t}$ converges to $\bm{1}\bm{\pi}^{\mathsf{T}}$. Because of (1), we obtain $\lim_{t\to\infty}\bm{x}(t)=\bm{1}\bm{\pi}^{\mathsf{T}}\bm{x}(0)$. We also use the diagonal matrix $\varPi:=\operatorname{diag}(\bm{\pi})$.
@@ 2.2 Performance measure of linear consensus algorithm
As a performance measure, we use the linear quadratic (LQ) cost [mainp]
@@
The cost (2) is obtained by evaluating $\mathrm{E}\left[\sum_{t\geq 0}\|\bm{x}(t)-\bm{x}(\infty)\|^{2}\right]$, where $\mathrm{E}[\cdot]$ means the expected value, under the assumption that $\bm{x}(0)$ is a random vector with covariance $\mathrm{E}\left[\bm{x}(0)\bm{x}(0)^{\mathsf{T}}\right]=I$ [aboutlqcost, mainp]. In addition, the cost (2) also appears in the noisy consensus [aboutlqcost, mainp, novelnoisy]. The noisy consensus repeats the update
@@
(3) $\frac{1}{n}\lim_{t\to\infty}\mathrm{E}\left[\|\tilde{\bm{e}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(t)}\|^{2}\right]$ $=$ $J(P).$
From the context of noisy consensus, we also use another variant [novelnoisy]
@@
which is obtained by substituting $\|\tilde{\bm{e}}\|^{2}$ in (3) with $\sum_{i}\pi_{i}\tilde{e}_{i}^{2}$. Unlike (2), which sums up the errors uniformly, (4) sums up the errors with weights according to $\bm{\pi}$.
@@ 2.3 Effective resistance
We consider a voltage vector of the nodes $\bm{v}$, where $v_{a}$ represents the voltage of the node $a$, and a current vector $\bm{i}$, where $i_{a}$ represents the current flowing out of (or into when negative) the node $a$. By Ohm’s law, the current from $a$ to $b$ through an edge $\{a,b\}$ is $C_{ab}(v_{a}-v_{b})$. Therefore, by Kirchhoff’s law, $i_{a}$ is determined by
@@
Because the current injection and extraction are balanced, we can assume $\bm{i}^{\mathsf{T}}\bm{1}=0$. Under this condition, the following lemma is well known fact (e.g., Lemma 6.12 in [FB-LNS]).
@@ 2.4 The relationships between linear consensus algorithm and effective resistance
The proof is analogous to the argument of Section 3.1.2 in [mainp].
Let $\Phi(P):=\Phi_{n}(P)=n\varPi P$, where $n$ denotes the number of rows of $P$. Then, we can see a clear relationship between the random walk with the transition matrix $P$ and some properties in the electrical network with conductance matrix $\Phi(P)$ such as the effective resistance, the voltage and the current (see [randomwalks]). In the analogy of random walk, a matrix called Green matrix [mainp] plays an important role. The Green matrix of consensus matrix $P$ is defined as
@@
This matrix is also called the fundamental matrix [revmarkov]. For a consensus matrix $P$, this matrix is well defined (see Chapter 11.5 in [greenconv]).
@@
Because $P^{t}-\bm{1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\pi}^{\mathsf{T}}}=(P-\bm{1}\bm{\pi})^{t}$ for $t\geq 1$, (7) can be reformulated as
@@
A fundamental matrix often refers not to $G(P)$, but to the right-hand side of (8) or (9) (e.g. [greenconv, novelnoisy]).
@@
which can be found in [mainp]. Such properties in this section are used to relate $J(P)$ and the effective resistance of the graph. In particular, $G(P)$ plays an important role in evaluating $J(P)$.
@@ 3 Performance analysis of linear consensus algorithm using effective resistance
In this section, we give upper and lower bounds for $J(P)$ using the concept of effective resistance. A bound for $J(P)$ has been known for reversible $P$ [mainp, novelnoisy], but not for nonreversible cases. To provide bounds for nonreversible $P$, we use a reversible matrix $P^{*}P$, where $P^{*}:=\varPi^{-1}P^{\mathsf{T}}\varPi$. Our result is a generalization of [mainp].
@@ 3.1 Properties of $P^{*}P$
We show some properties of $P^{*}P$. For every consensus matrix $P$, $P^{*}P$ is a reversible consensus matrix with the same invariant measure as $P$, so it is called a multiple reversiblization of $P$ [reversiblization]. When $P$ is a reversible consensus matrix, we have $P^{*}=\varPi^{-1}(\varPi P)^{\mathsf{T}}=\varPi^{-1}(\varPi P)=P$ and $P^{*}P=P^{2}$.
@@
Let $Q:=\varPi^{1/2}P\varPi^{-1/2}$. Then, $\operatorname{tr}\left((P^{*})^{t}P^{t}\right)=\operatorname{tr}\left((Q^{t})^{\mathsf{T}}Q^{t}\right)=\sum_{i}\sigma_{i}^{2}(Q^{t})$, where $\sigma_{i}(Q)$ is the $i$-th singular value of $Q$. Moreover, $\operatorname{tr}\left((P^{*}P)^{t}\right)=\operatorname{tr}\left((Q^{\mathsf{T}}Q)^{t}\right)=\sum_{i}\sigma_{i}^{2t}(Q)$, because $(Q^{\mathsf{T}}Q)^{t}$ is symmetric. Using $\sigma_{i}(Q^{t})\leq\sigma_{i}^{t}(Q)$, as shown in 9.H.2.a of [traceineq], we obtain (10).
The equality of (10) holds not only for reversible matrices. For example, normal consensus matrices satisfy the property $P^{*}P=PP^{*}$. The class of normal consensus matrices includes Cayley graphs, and $J(P)$ on Cayley graphs have been studied (e.g. [cayley]).
@@
Because $P$ is normal, $(I-P)(I-P^{\mathsf{T}})=(I-P^{\mathsf{T}})(I-P)$. Multiplying $\bm{1}$ from the right, we obtain $(I-P)(I-P^{\mathsf{T}})\bm{1}=\bm{0}$, which means $(I-P^{\mathsf{T}})\bm{1}\in\ker(I-P)$. Therefore, there exists some real value $a$ that satisfies $(I-P^{\mathsf{T}})\bm{1}=a\bm{1}$, because $\ker(I-P)=\operatorname{span}\{\bm{1}\}$. Then, we multiply $\bm{1}^{\mathsf{T}}$ from the left and obtain $\bm{0}=an$. This means that $a=0$ and $(I-P^{\mathsf{T}})\bm{1}=\bm{0}$. Thus, $P$ is doubly-stochastic and the invariant measure $\bm{\pi}$ is $\frac{1}{n}\bm{1}$. Therefore, we obtain $P^{*}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}P^{\mathsf{T}}}$ and $P^{*}P=P^{\mathsf{T}}P=PP^{\mathsf{T}}=PP^{*}$.